# Interior curvature estimate for curvature quotient equations on convex hypersurfaces

### v1

@@ 1. Introduction
Inspired by Lu’s [13], we rederive Lu’s Jacobi inequality specifically tailored for Hessian quotient equations. Unlike the integral method employed in Lu’s [13] which refer to hankar and Yuan’s [20], our proof takes a pointwise approach, thereby circumventing the challenges posed by the use of the Legendre transform on Riemannian manifolds. Building on the auxiliary function and the case-by-case method introduced by Guan and Qiu [6], we examine their applicability in the setting of curvature quotient equations and obtain a result that is analogous in spirit to theirs. While many previous studies have focused on low-dimensional cases ($n=3$ or $n=4$), the method considered here appears to be applicable to general dimensions $n\geq 3$, potentially contributing to the broader study of interior estimates for curvature equations.
@@ 3. Jacobi inequality
(4.23) $F^{ij}P_{ij}$ $\geq-\frac{4C}{\lambda_{n}}-C\left(\frac{1}{\rho}+\beta+\alpha\right)+\frac{Fh_{11}}{\log h_{11}}+\frac{C\alpha}{\lambda_{n}}.$
Since $\epsilon$ is fixed, choosing $\alpha\geq 4$, we obtain:
@@
Here, $j$ is fixed throughout this expression, so $F^{jj}h_{11j}$ does not represent a summation.
@@
Analyze the coefficient in right hand side of this inequality:
@@
Then we choose $\alpha\geq\beta\geq 4$ and obtain:
@@
② If $|(e_{j},E_{n+1})|\geq A>0$, where $A\in(0,1]$ is a positive constant, then since $|b_{j}|<C(|M|_{C^{1}(B_{1})},n)$, by firstly choosing $\beta\geq 2$ and then $\alpha\geq max\{4,\frac{2\beta(\|M\|_{C^{1}(B_{r})}+d)}{A}\}$ we have:
@@
$\alpha\geq max\{4,\frac{2\beta(\|M\|_{C^{1}(B_{r})}+d)}{A},\frac{8d}{A^{2}},\frac{2\sqrt{d\beta(\|M\|_{C^{1}(B_{r})}+d)}}{A}\}$
@@ Acknowledgements
This work was carried out while the author was a Master’s student at McMaster University, and constitutes the author’s Master’s thesis. The current version also corrects certain errors present in the original thesis manuscript. The author would like to thank Professor Siyuan Lu for helpful discussions and for bringing many related works to our attention. He is also deeply grateful to Bin Wang for numerous important discussions and for his careful suggestions regarding the correctness of this work. In particular, the author sincerely thanks Dr. Guohuan Qiu for kindly reviewing the manuscript with great care.

### v2

@@ 1. Introduction
Inspired by Lu’s work [13], we revisit the Jacobi inequality on Riemannian manifolds. In contrast to the integral method developed by Lu, which draws on the framework of Shankar and Yuan [20], our approach is pointwise, thus avoiding the technical difficulties of applying the Legendre transform on Riemannian manifolds. Building on the auxiliary function technique of Guan and Qiu [6], we develop a new case-by-case analysis tailored to the structure of Hessian quotient equations and obtain a structurally similar result. The cases $n=3$ or $4$ have been well studied, while higher dimensions remain less understood. Our method applies to general $n\geq 3$ and may offer a modest contribution.
@@ 3. Jacobi inequality
(4.23) $F^{ij}P_{ij}$ $\geq-\frac{C}{\lambda_{n}}-C\left(1+\beta+\alpha\right)+\frac{Fh_{11}}{\log h_{11}}+\frac{C\alpha}{\lambda_{n}}.$
Choosing $\alpha$ sufficiently large, we obtain:
@@
Here, $j$ is fixed throughout this expression, so $F^{jj}h^{2}_{11j}$ does not represent a summation.
@@
Assume first that $\alpha$ is sufficiently large so that Case 1 applies. We then analyze the coefficient on the right-hand side of the inequality:
@@
Then we choose $\beta\geq 2$ to obtain:
@@
② If $|(e_{j},E_{n+1})|\geq A>0$, where $A\in(0,1]$ is a positive constant, then since $|b_{j}|<C(|M|_{C^{1}(B_{1})},n)$, by firstly choosing $\beta\geq 2$ and then $\alpha\geq\frac{2\beta(\|M\|_{C^{1}(B_{r})}+d)}{A}$ we have:
@@
$\alpha\geq max\{\frac{2\beta(\|M\|_{C^{1}(B_{r})}+d)}{A},\frac{8d}{A^{2}},\frac{2\sqrt{d\beta(\|M\|_{C^{1}(B_{r})}+d)}}{A}\}$
@@ Acknowledgements
This work was carried out while the author was a Master’s student at McMaster University, and constitutes the author’s Master’s thesis. The current version also corrects certain errors present in the original thesis manuscript. The author would like to thank Professor Siyuan Lu for helpful discussions and for bringing many related works to our attention. He is also deeply grateful to Bin Wang for numerous important discussions and for his careful suggestions regarding the correctness of this work. In particular, the author sincerely thanks Professor Guohuan Qiu for kindly reviewing the manuscript with great care.